# ICD-LM: Configuring Vision-Language In-Context Demonstrations by Language Modeling

### v2

@@ 4 Experiments
4.1 Tasks and Datasets
Our approach is evaluated on MS-COCO for Image Captioning (IC) and VQAV2 for Visual Question Answering (VQA). For each corresponding dataset, we use the train split to construct the $\mathcal{D}_{\mathcal{M}}$ for training ICD-LM and use the validation split to evaluate the performance of ICD configurations generated by ICD-LM.
MS-COCO [16]: MS-COCO is widely used in IC, which is divided into $118,287$ training, $5,000$ validation, and $5,000$ test image-caption pairs. Notably, each training image is associated with five distinct human-annotated captions.
VQAV2 [10]: VQAV2 emphasizes open-ended VQA tasks, which encompasses $4,437,570$ question-answer pairs in its training split, supplemented by an additional $2,143,540$ pairs in the validation split.
4.2 Implementation Details
Constructing $\mathcal{D}_{\mathcal{M}}$: We construct 2-shot $\mathcal{D}_{\mathcal{M}}$ for main experiments. We select $5000$ samples to get the anchor set $\mathcal{A}$. For each anchor sample, we randomly choose $64$ samples to build the sub-supporting set $\mathcal{D}^{\bm{a}}_{S}$. The beam size for these processes is 5. More details are in Appendix.
Training ICD-LM: Different training strategies are employed for IC and VQA when training ICD-LM. In IC, the weight of CLIP model will be frozen, and an MLP adapter is introduced to its output. While, for VQA, the CLIP encoder remains trainable, and no adapter is appended. Both the ICD-LM for VQA and IC use 2-layer Transformer decoder blocks [34] or LSTM [27] as the architecture of LM. The training phase leverages the AdamW optimizer [19] and a cosine learning rate scheduler. We set the learning rate to $1\times 10^{-4}$ and the batch size to 128. We train our ICD-LM for 20 epochs. All training processes are carried out with mixed precision and 2 RTX3090 GPUs.
Implementing ICL: To implement ICL, we use OpenFlamingoV2-9B [2] as our LVLM. We use beam search during inference where the beam size is set to 3. Besides, we set the maximum number of generated tokens as 20 in IC and 5 in VQA. All experiments are deployed on an RTX 3090, operating in BF16 mode. The prompt structures of IC and VQA are shown in Table  1.
Table 1: Prompt Formats for IC and VQA tasks with placeholders.
Task Prompt Format
VQA <image>Question:<Q> Short answer:<A>
Caption <image>Output:<X>
@@ 4.3.1 Comparison Methods
We compare ICD-LM to 4 ICD selection strategies:
@@
Similarity-based Retrieval methods: These methods form $\mathcal{S}^{k}$ by computing the similarity between the query input $\mathbf{x}^{\prime}$ and ICDs in $\mathcal{D}_{S}$. Specifically, we use CLIP to extract the features of text or image and use cosine similarity to measure the similarity between the extracted features. We follow  [18], sorting examples in ascending order by their similarity to the query input, so the rightmost demonstration is the closest example. (1). Similarity-based Image-Image Retrieval (SIIR): We select $k$ ICDs from $\mathcal{D}_{S}$ with highest image similarity to the query image. (2). Similarity-based Text-Text Retrieval (STTR): We select $k$ ICDs from $\mathcal{D}_{S}$ that have most similar texts to the text of the query. For VQA, with question as input, we can use STTR to form $\mathcal{S}^{k}$. While for IC, the query input is only an image without any text input, thus STTR is infeasible. (3). Similarity-based Image-Text Retrieval (SITR): Taking advantage of the cross-modal retrieval capability of CLIP, we can compute the similarity between different modalities. Specifically, we compute the similarity between query image and all text of $d_{i}\in\mathcal{D}_{S}$ and select ICDs whose texts have the top-$k$ similarities with the query image. The text input varies per task, e.g., it is the caption for IC and the question for VQA.
@@ 4.3.2 Main Result
Overall, ICD-LM achieves the best performance on most cases compared to other methods. Notably, ICD-LM consistently outperforms other methods for the Avg:1$\sim$2. Specifically, in IC, ICD-LM surpasses the best similarity-based retrieval method, SIIR, by 6.03 in CIDEr score (84.32 vs. 78.29). Similarly, for VQA, ICD-LM exceeds the performance of the best similarity-based retrieval method, STTR, by achieving an accuracy 3.07 higher (48.75 vs. 45.68). The results indicate that ICD-LM performs well in generating 2-shot ICD configurations.
Moreover, we find that ICD-LM has remarkable extrapolation abilities. Regarding the Avg:3$\sim$8 metric, ICD-LM maintains the top performance in both IC and VQA. Specially, compared with SIIR, ICD-LM achieves 0.8 higher CIDEr in IC (96.52 vs. 95.72) and 0.75 higher accuracy in VQA (52.59 vs. 51.84). Thus, even with a dataset consisting of only 2 shots, ICD-LM is still able to generate longer high-quality ICD configurations. These experiment results indicate that ICD-LM can effectively generate optimal ICD configurations, thereby assisting LVLM in predictions.
Furthermore, we observe the ICD configurations generated by ICD-LM are more robust. We can observe the SITR underperforms when compared to RS in both IC and VQA.
For STTR in VQA, it exhibits a performance decline compared to RS when the number of shots is relatively large. For instance, it lags behind RS by 0.81 in the 8-shot configuration. These results highlight that not all similarity-based retrieval methods are effective in different multimodal tasks.
In Fig. 4, we visualize several ICD configurations and observe that, when the context of the ICDs is similar to query sample, OpenFlamingo tends to excessively rely on text information of ICDs when predicting labels. For example, in Fig. 4 (a) and (b), all the ICD questions are yes-or-no type, which causes LVLMs to output “no” even when faced with a “What” type question of the query sample. While for IC in Fig. 4 (c) and (d), SITR incorrectly outputs “London” even though the location is not explicitly indicated in the query image, while SIIR even leads LVLMs to directly copy the text from the ICD. This behavior, identified as Short-cut Inference, has been previously highlighted in a study [41]. Conversely, ICD-LM generates more diverse ICD configurations, thereby preventing misleading inferences.
Table 2: Results of diverse ICL methods on IC and VQA.

### v4

@@ 4 Experiments
4.1 Datasets and implementation details
Our approach is evaluated on MS-COCO [56] for Image Captioning (IC) and VQAV2 [60] for Visual Question Answering (VQA). For each corresponding dataset, we use the train split to construct the $\mathcal{D}_{\mathcal{M}}$ and use the validation split to evaluate the performance of ICD configurations generated by Lever-LM. More details are given in Appendix A.
To get $\mathcal{D}_{\mathcal{M}}$, we select $5000$ samples to get the anchor set $\mathcal{A}$. For each anchor sample, we randomly choose $64$ samples to build the sub-supporting set $\mathcal{D}^{\bm{a}}_{S}$. The beam size for these processes is 5. To train Lever-LM, different strategies are employed for IC and VQA. In IC, the weight of CLIP model will be frozen, and an MLP adapter is introduced to its output. While, for VQA, the CLIP encoder remains trainable, and no adapter is appended. The training phase leverages the AdamW optimizer [61] and a cosine learning rate scheduler. We set the learning rate to $1\times 10^{-4}$ and the batch size to 128. We train our Lever-LM for 20 epochs. To implement ICL, we use OpenFlamingoV2-9B [34] and IDEFICS-9B [14] as our LVLMs. We use beam search during inference where the beam size is set to 3. Besides, we set the maximum number of generated tokens as 20 in IC and 5 in VQA.
@@ 4.2.1 Comparison Methods
We compare Lever-LM with 4 ICD selection strategies:
@@
Similarity-based Retrieval methods: To date, only a few studies focus on configuring ICD sequence for solving VL tasks [24, 25], where both studies show that, despite their simplicity, similarity-based retrieval methods are effective for selecting ICDs. We therefore consider these strategies as current SOTA benchmarks to assess our effectiveness.  [^3] Note that we use different experiment settings from [24] and more details are given in Appendix A.4. They form $\mathcal{S}^{k}$ by computing the cosine similarity between the query input $\mathbf{x}^{\prime}$ and ICDs in $\mathcal{D}_{S}$ where CLIP is used to extract features. We follow [25] to sort examples in ascending order by their similarity to the query input, so the rightmost demonstration is the closest example. Similarity-based methods contain three variants: (1). Similarity-based Image-Image Retrieval (SIIR): We select $k$ ICDs from $\mathcal{D}_{S}$ with highest image similarity to the query image. (2). Similarity-based Text-Text Retrieval (STTR): We select $k$ ICDs from $\mathcal{D}_{S}$ with highest text similarity to the query text. This method is only applicable to VQA where question is used as text and not infeasible for IC. (3). Similarity-based Image-Text Retrieval (SITR): We compute the similarity between query image and all text of $d_{i}\in\mathcal{D}_{S}$ and select ICDs whose texts have the top-$k$ similarities with the query image. For IC/VQA, we use caption/question for IC/VQA.
@@ 4.2.2 Main Result
The results for various ICD selection strategies are shown in Table 1 for IC and VQA. For Lever-LM, it is trained by $\mathcal{D}_{\mathcal{M}}$ whose ICD length is set to 2. Due to increased inference time with more shots, we do not test the inference results for 5- and 7-shots. The table shows the length interpolation and extrapolation ability of Lever-LM. Interpolation refers to performance with ICDs shorter than those in the training set $\mathcal{D}_{\mathcal{M}}$, which contains only 2-shot ICDs, and is denoted as "Avg:1$\sim$2". Extrapolation pertains to performance with ICDs longer than those in $\mathcal{D}_{\mathcal{M}}$, represented as "Avg:3$\sim$8". The notation "Avg:1$\sim$8" indicates overall performance across 1 to 8 shots. Future analysis will focus on comparing these averages to minimize potential bias across methods.
Overall, Lever-LM achieves the best performance on most cases compared to other methods on both LVLMs. Notably, Lever-LM excels in Avg:1$\sim$2. Specifically, in VQA, Lever-LM surpasses the best performing SIIR method by 3.07 (48.75 vs. 45.68) and 0.57 (53.65 vs. 53.08) in accuracy on the IDEFICS and OpenFlamingo models, respectively. In IC, Lever-LM outperforms the best baseline, SIIR, by 6.03 (84.32 vs. 78.29) CIDEr on OpenFlamingo. Similarly, for IDEFICS, Lever-LM achieves a higher CIDEr of 3.2 (89.57 vs. 86.37) compared to the best baseline, RS.
Moreover, Lever-LM has remarkable extrapolation abilities. Regarding Avg:3$\sim$8, Lever-LM maintains the top performance in both IC and VQA. Specially, on OpenFlamingo, Lever-LM outperforms SIIR with a 0.8 higher CIDEr in IC (96.52 vs. 95.72) and a 0.75 greater accuracy in VQA (52.59 vs. 51.84). Meanwhile, on IDEFICS, compared with RS, Lever-LM achieves a 2.93 higher CIDEr in IC (105.79 vs. 102.86) and 0.49 higher accuracy in VQA (54.84 vs. 54.35). These results indicate that Lever-LM can effectively identify and utilize internal statistical patterns to compose longer, high-quality ICD sequences, even from a dataset comprising only two shots.
When we delve deeper into the results in Table 1, we find that the relative performance of similarity-based methods and RS varies by LVLM and task. For example, for IC, SIIR outperforms RS ( Avg:1$\sim$8: 89.91 vs. 88.48) when OpenFlamingo is used to implement ICL while SIIR significantly lags behind RS (Avg:1$\sim$8: 88.19 vs. 97.36) when IDEFICS is used. Also, for VQA, STTR is comparable to RS (Avg:1$\sim$8: 47.98 vs. 47.94) on OpenFlamingo while STTR is defeated by RS (Avg:1$\sim$8: 49.75 vs. 53.54) when IDEFICS is used. These performance fluctuations demonstrates the instability of these heuristic-based methods. However, Lever-LM does not have such serious fluctuations where it outperforms both RS and similarity-based retrieval methods across various LVLMs and tasks on average. Such observations also suggest that Lever-LM may capture the stable statistic patterns between ICDs.
Besides the above-mentioned advantages, Fig. 3 shows that Lever-LM is more robust to the Short-cut Inference brought by using similarity-based retrieval methods [24, 25]. For example, in (a) and (b), all the ICD questions are yes-or-no type, which causes LVLMs to output “no” even when faced with a “What” type question of the query sample. For IC in (c) and (d), SITR incorrectly outputs “London” even though the location is not explicitly indicated in the query image, while SIIR even leads LVLMs to directly copy the text from the ICD. Conversely, Lever-LM generates more diverse ICD configurations, thereby preventing misleading inferences.
Table 1: Results of diverse ICL methods on IC and VQA, where “OF” and “IDE” denote OpenFlamingo and IDEFICS, respectively. Lever-LM is trained by $\mathcal{D}_{\mathcal{M}}$ whose ICD length is set to 2.
Interpolation Extrapolation Avg:1$\sim$8
Shot 1 Shot 2 Avg:1$\sim$2 Shot 3 Shot 4 Shot 6 Shot 8 Avg:3$\sim$8
OF IC RS 73.32 82.95 78.14 87.72 93.65 95.81 97.42 93.65 88.48
SITR 66.05 77.69 71.87 83.46 85.05 89.84 93.57 87.98 82.61
SIIR 71.71 84.87 78.29 90.83 93.22 97.80 101.01 95.72 89.91
Lever-LM 80.02 88.63 84.32 93.41 96.06 97.26 99.35 96.52 92.45
OF VQA RS 41.97 45.92 43.95 48.17 48.95 51.18 51.44 49.94 47.94
SITR 40.17 43.58 41.88 46.03 47.5 49.72 50.75 48.50 46.29
SIIR 43.31 47.46 45.39 49.85 50.68 53.23 53.58 51.84 49.69
STTR 44.6 46.75 45.68 47.92 49.05 50.06 49.47 49.13 47.98
Lever-LM 46.66 50.83 48.75 51.91 52.15 53.29 53.01 52.59 51.31
IDE IC RS 76.44 96.31 86.37 100.80 101.82 103.64 105.18 102.86 97.36
SITR 62.08 75.50 68.79 82.57 86.64 90.34 92.88 88.11 81.67
SIIR 66.61 83.31 74.96 89.43 93.02 97.07 99.70 94.81 88.19
Lever-LM 78.70 100.45 89.57 104.52 104.86 106.44 107.33 105.79 100.38
IDE VQA RS 52.40 53.21 52.81 53.47 53.70 54.00 54.48 53.91 53.54
SITR 51.52 51.72 51.62 52.55 52.59 52.83 50.49 52.12 51.95
SIIR 52.87 53.28 53.08 53.77 53.92 54.81 54.88 54.35 53.92
STTR 48.13 49.32 48.73 49.43 50.11 50.58 50.93 50.26 49.75
Lever-LM 53.31 53.98 53.65 54.39 54.58 55.09 55.3 54.84 54.44
4.3 Ablation Studies
We use ablation studies to explore the effects of diverse settings on our approach, including (1) diverse $\mathcal{D}_{\mathcal{M}}$ configurations; (2) diverse scorers $I_{\mathcal{M}}$ in Eq. (3); (3) diverse LM structures; (4) $\mathcal{D}_{\mathcal{M}}$ with 4-shot ICDs; and (5) randomly ordering the ICD sequences generated by Lever-LM.
Diverse $\mathcal{D}_{\mathcal{M}}$ Configurations. We generate 2-shot $\mathcal{D}_{\mathcal{M}}$ by different settings to investigate the corresponding effects. Three factors are ablated: beam size $b$; the number $n$ of samples in $\mathcal{A}$; and the sampling method of $\mathcal{D}_{S}^{\bm{a}}$, including 3 methods: Random: Selecting randomly from $\mathcal{D}_{S}$; Similar Text (Sim-T): Selecting the highest textual similarity sample with anchor sample $\bm{a}$ from $\mathcal{D}_{S}$; Similar Image (Sim-I): Selecting the highest visual similarity sample with anchor sample $\bm{a}$ from $\mathcal{D}_{S}$.
Table 2 (3) $\sim$ (11) shows the results for different $\mathcal{D}_{\mathcal{M}}$ configurations on IC and VQA. We find that Lever-LM can consistently improve the performance compared with the baseline RS  [^4] The RS means Random Sample ICD retrieval method which is mentioned in Section 4.2.1. in Avg:1$\sim$2. As for length extrapolation capability, only Sim-T gets a lower score than RS. These comparisons confirm the robustness of our method.
Table 2 (3) $\sim$ (5) show that appropriately increasing the beam size $b$ can improve performance. Specifically, as $b$ increases from $1$ to $5$, in Avg:1$\sim$8, CIDEr/accuracy increases by 2.79/1.73 for IC/VQA. This suggests that a diverse $\mathcal{D}_{\mathcal{M}}$ encompasses a broader range of high-quality ICD configurations, which can help train a better Lever-LM. However, an excessively large $b$ can negatively impact performance. For instance, in VQA, the accuracy of $b=10$ decays 0.12 than $b=5$ in Avg:1$\sim$8. We hypothesize this drop in performance is due to the introduction of lower-scoring ICD sequences with a large beam size, potentially misleading Lever-LM during training.
Table 2 (6) $\sim$ (8) show that using more anchor samples can improve the interpolation performance in both IC and VQA, e.g., when $n$ increases from 1000 to 5000, the Avg:1$\sim$2 CIDEr/accuracy of IC/VQA increases from 83.94/45.39 to 84.32/ 48.75. However, we find that on IC, although the interpolation performance increases when $n$ changes from 3000 to 5000, the extrapolation performance decays, e.g., Avg:3$\sim$8 decreases from 97.60 to 96.52. One possible reason is that using Eq. (3) to build $D_{\mathcal{M}}$ may introduce certain in-domain bias which is beneficial for interpolation while detrimental for extrapolation on IC.
For different sample methods of constructing the $\mathcal{D}_{S}^{a}$ in table 2 (9) $\sim$ (11), we find Random is the best in both IC and VQA. We suppose this is because selecting similar ICDs with the anchor sample from $\mathcal{D}_{S}$ will damage the diversity. Previous study [62] in NLP validates that the diversity of the ICD sequences will also help improve the performance of LLMs.
Table 2: Results of diverse ablation studies on IC and VQA.
IC VQA
Avg:1$\sim$2 Avg:3$\sim$8 Avg:1$\sim$8 Avg:1$\sim$2 Avg:3$\sim$8 Avg:1$\sim$8
(1) RS 78.14 93.65 88.48 43.95 49.94 47.94
(2) Lever-LM 84.32 96.52 92.45 48.75 52.59 51.31
(3) $b=1$ 79.91 94.53 89.66 46.47 51.13 49.58
(4) $b=5$ 84.32 96.52 92.45 48.75 52.59 51.31
(5) $b=10$ 84.96 97.12 93.06 48.58 52.49 51.19
(6) $n=1000$ 83.94 96.74 92.48 45.39 50.44 48.76
(7) $n=3000$ 84.13 97.60 93.11 47.56 51.11 49.93
(8) $n=5000$ 84.32 96.52 92.45 48.75 52.59 51.31
(9) Sim-I 81.96 96.11 91.40 47.10 51.79 50.23
(10) Sim-T 81.22 87.66 85.52 45.38 49.55 48.16
(11) Random 84.32 96.52 92.45 48.75 52.59 51.31
(12) CIDEr Scorer 87.93 93.52 91.65 - - -
(13) Lever-LM LSTM 83.93 96.21 92.12 46.60 50.68 49.32
(14) Golden-1 81.78 97.44 92.22 47.78 52.51 50.93
(15) Golden-2 91.20 99.63 96.82 45.32 49.05 47.80
Diverse scorers $I_{\mathcal{M}}$ for evaluating ICD sequences. To evaluate the quality of ICD configurations, we can use task-specific metrics as $I_{\mathcal{M}}$ to build $\mathcal{D}_{\mathcal{M}}$, such as CIDEr in IC. Table 2 (2) and (12) compare the results between using prediction confidence Eq. 3 (2) and CIDEr (12) as $I_{\mathcal{M}}$. We find that using CIDEr achieves 3.61 higher than Confidence in Avg:1$\sim$2, suggesting that CIDEr can assign a more accurate and reasonable score for ICD configurations. However, the length extrapolation capability decreases obviously, which is 3.0 lower than Confidence in Avg:3$\sim$8, validating the robustness of Confidence scorer. Moreover, it will cost more time to construct $\mathcal{D}_{\mathcal{M}}$ by task-specific metric is used, e.g., CIDEr costs approximately 10 times of Confidence when constructing $\mathcal{D}_{\mathcal{M}}$.
Figure 3: Visualizations of diverse ICDs configurations, where the first and the last ICDs are given due to space limitation. We can find that Lever-LM use more diverse ICDs and thus not lead to short-cut inference.
Diverse LM Structures. Table 2 (13) shows the results of using LSTM [63] as Lever-LM, we find that this still achieves excellent performance. For example, in IC, the overall performance improves by 3.64 (92.12 vs. 88.48) compared to the RS baseline, while in VQA, it is improved by 1.38 (49.32 vs. 47.94). However, due to the weak representation learning capability of LSTM, its performance is lower than Transformer,e.g., the scores decrease by 0.33/1.99 in IC/VQA, respectively. Overall, these results suggest that effective ICD configurations contain internal statistic patterns which can be captured by different temporal learner.
Table 3: Results of Lever-LM with 4-shot $\mathcal{D}_{\mathcal{M}}$ on IC and VQA.
Model IC VQA
Avg:1$\sim$4 Avg:6$\sim$8 Avg:1$\sim$8 Avg:1$\sim$4 Avg:6$\sim$8 Avg:1$\sim$8
RS 84.41 96.62 88.48 46.25 51.31 47.94
SITR 78.06 91.71 82.61 44.32 50.24 46.29
SIIR 85.16 99.40 89.91 47.83 53.41 49.69
STTR - - - 47.08 49.77 47.98
Lever-LM(4-shot $\mathcal{D}_{\mathcal{M}}$) 87.35 97.96 90.88 48.56 52.68 49.93
Lever-LM(2-shot $\mathcal{D}_{\mathcal{M}}$) 89.53 98.30 92.45 50.39 53.15 51.31
Table 4: Results of Random Order of Lever-LM generated ICDs.
2-Shot $\mathcal{D}_{\mathcal{M}}$ 4-Shot $\mathcal{D}_{\mathcal{M}}$
VQA IC VQA IC
Original 50.83 88.63 51.12 85.97
Random Order 50.42 88.56 50.63 85.77
Golden ICD Sequence. In experiments, we find that using a high learning rate and not freezing the CLIP model may make Lever-LM converge to a specific solution that for any query input, the ICD configuration is fixed while still return good ICL performance. For example, in IC, the best version, Golden-2, can outperform the non-Fixed case (Table 2 (2)) by 4.37 points in Avg:1$\sim$8. Such improvement suggest that if we do not have enough computation burdens to configure diverse ICD sequence for each query, we can preserve one Golden ICD Sequence for the latter usage. However, we also find that the performance of Golden ICD Sequence fluctuates significantly, e.g., Golden-2 is poorer than RS in VQA in Avg:1$\sim$8. This also points out a new direction to study how to get more stable Golden ICD Sequence.
Longer Few-shot $\mathcal{D}_{\mathcal{M}}$. We further explore the performance of Lever-LM using a 4-shot $\mathcal{D}_{\mathcal{M}}$, as presented in Table 3. It is evident that Lever-LM continues to outperform other retrieval-based methods in Avg:1$\sim$8 metric. However, we observe a notable performance reduction when Lever-LM is trained with the 4-shot $\mathcal{D}_{\mathcal{M}}$ compared to the 2-shot $\mathcal{D}_{\mathcal{M}}$. Specifically, for IC, there is a performance decrease of approximately 1.57 in the Avg:1$\sim$8 metric when using the 4-shot $\mathcal{D}_{\mathcal{M}}$ to train Lever-LM. One possible reason is that when constructing $\mathcal{D}_{\mathcal{M}}$, some approximation operations are applied where sub-optimal ICD sequences are got. Then parts of statistic patterns may be salient is $\mathcal{D}_{\mathcal{M}}$ that are more easily captured by Lever-LM where using longer ICD sequences may further encourage Lever-LM to capture such patterns, and thus causing less effective ICD sequences than the shorter $\mathcal{D}_{\mathcal{M}}$. This points out a new direction to study how to build $\mathcal{D}_{\mathcal{M}}$ with longer and more robust ICD sequences for better training Lever-LM.
Random Order ICD sequence. To validate that whether Lever-LM captures effective ICD orders, we randomly rearrange the ICD sequences generated by Lever-LM trained with 2-shot and 4-shot $\mathcal{D}_{\mathcal{M}}$ and then evaluate the performance the 2-shot and 4-shot ICD configurations, respectively, in Table 4. It is evident that the original order of ICDs generated by Lever-LM attains the highest score in both VQA and IC, validating that Lever-LM can learn how to order the ICDs.
5 Conclusion
After observing that configuring an ICD sequence is a mirror process of composing a sentence, we assume effective ICDs may contain statistic patterns that can be captured by temporal learner. Then we use a tiny LM named as Lever-LM to capture such patterns for configuring ICDs to lever LVLMs. To achieve this, we construct a dataset containing effective ICD sequences to train this Lever-LM. After training, we validate the effectiveness of Lever-LM by comparing it with similarity-based retrieval methods and find that Lever-LM can capture the statistic patterns between ICDs. Extensive ablations are deployed to discover which factors and why they will affect the results, which also pointing out a few future research directions.
Acknowledgement
This work is supported by the National Science Foundation of China (62206048), the Natural Science Foundation of Jiangsu Province (BK20220819), the Young Elite Scientists Sponsorship Program of Jiangsu Association for Science and Technology (Tj-2022-027), and the Fundamental Research Funds for the Central Universities (2242024k30035). This research work is also supported by the Big Data Computing Center of Southeast University.
Appendix A Implementation Details.
A.1 Lever-LM Hyperparameters
We provide additional implementation details for each experiment in Table 5. For all experiments, the batch size is set to 64, the warmup steps to $5\%$ of total training steps, the scheduler as a cosine scheduler, and the optimizer as AdamW [61]. All experiments are deployed on an RTX 3090. All training processes are carried out with mixed precision and 2 RTX3090 GPUs. While for LVLM ICL experiments, we use BF16 mode.
Table 5: Different settings of Lever-LM experiments, where the $n$ is the number of anchor samples in $\mathcal{A}$, $b$ is the beam size, and $l$ is the length of ICD configurations.
Training Parameters $\mathcal{D}_{\mathcal{M}}$ Parameters Avg:1$\sim$8
lr weight decay epoch Freeze Adapter $n$ $b$ $l$
@@ A.1 Lever-LM Hyperparameters
[…]
A.2 Datasets
MS-COCO [56] is widely used in IC, which is divided into $118,287$ training, $5,000$ validation, and $5,000$ test image-caption pairs. Notably, each training image is associated with five distinct human-annotated captions.
VQAV2 [60] emphasizes open-ended VQA tasks, which encompasses $4,437,570$ question-answer pairs in its training split, supplemented by an additional $2,143,540$ pairs in the validation split.
A.3 Prompt Template
For different LVLMs and tasks, the input format also varies. We followed the prompt templates provided by OpenFlamingo and IDEFICS for our experiments. We show the prompt templates in table 6. In the VQA tests of IDEFICS, an additional instruction needs to be added at the beginning of the input. Therefore, our final input format of k-shot is: [instruction] + [ICD1 prompt] + [ICD2 prompt] + … [ICDk prompt] + [Query prompt].
Table 6: Prompt Formats for IC and VQA tasks with placeholders.
LVLM Task Prompt Template Instruction
IDEFICS IC Caption:<X> -
VQA Question:<Q> Short answer:<A> provide an answer to the question. Use the image to answer.
OpenFlamingo IC Output:<X> -
VQA Question:<Q> Short answer:<A> -
A.4 Similarity-based retrieval method
In our study, we adopt the methodology proposed by  [24] as a baseline, yet we incorporate several distinct experimental settings. Firstly, unlike their use of OpenFlamingoV1 [34] which employs LLaMA [5] as the underlying LLM, we utilize OpenFlamingoV2, based on the MPT LLM [37]. This version of OpenFlamingo leverages a more extensive dataset and a more advanced LLM, resulting in enhanced robustness and improved ICL capabilities.
Secondly, while they focus exclusively on ICL strategies in IC, which lacks textual input, our research extends to VQA. VQA involves textual queries; hence, we have adapted the STTR retrieval method for sourcing samples with similar textual prompts as ICDs, drawing inspiration from their approach.
Lastly, in the context of Image Captioning, our experiment utilizes the MSCOCO 2017 dataset, in contrast to their employment of the Karpathy split [64] of the MSCOCO 2014 dataset.
Appendix B Detail Experiments Result
In this section, we present detailed experimental results and ablation studies on OpenFlamingo and IDEFICS models. The results include CIDEr scores, CHAIR metrics, fine-grained analyses, model size comparisons, inference time comparisons, and CLIPScore evaluations.
B.1 Detail Ablation Studies Results On OpenFlamingo
In Table 7, Table 8, Table 9, Table 8, Table 10 and Table 11, we present all the detail ablation studies results on OpenFlamingo.
Table 7: Results of diverse $\mathcal{D}_{\mathcal{M}}$ configurations on IC and VQA.